OpenStudy (anonymous):

System of Equations, Solve for x 8x+2y=6 3x+2y=7

Parth (parthkohli):

Use the elimination method, that is, you subtract both equation 1 and 2 and you have one variable left.

OpenStudy (anonymous):

13/11 13/5 -1/5 1/5

OpenStudy (anonymous):

i think it would be B 13/5

Parth (parthkohli):

5x = -1 after subtracting both equations. Now, can you do it?

OpenStudy (anonymous):

so its 1/5

OpenStudy (anonymous):

8x+2y=6 3x+2y=7 subtract first equation to second equation 5x = -1 x= -1/5